# Compositional Zero-shot Learning via Progressive Language-based Observations

### v1

@@ 4.1 Experiment Settings
Datasets. We evaluate performance on three challenging benchmarks: MIT-States [12]: It comprises 53,753 natural images, with 115 states and 245 objects. In the closed-world setting, the search space includes 1,262 seen compositions and 300/400 unseen compositions for validation/testing. 2) UT-Zappos [29]: It consists of 50,025 images of shoes, with 16 states and 12 objects. In the closed-world experiments, we considered 83 seen compositions and 15/18 (validation/test) unseen compositions following the constraints defined in [33]. 3) C-GQA [39]: It contains 453 states and 870 objects, comprising a total of 39,298 images. The dataset is divided into 5,592 seen compositions for training, and 1,040/923 unseen compositions for validation/testing, respectively. In the open-world settings, these datasets contain 28,175, 192, and 278,362 compositions, respectively.
@@
Metrics. We followed the established CZSL evaluation protocol [25] and assessed all results using four metrics in both closed-world and open-world scenarios: 1) Seen (S): This metric measures the accuracy specifically for seen compositions. 2) Unseen (U): It evaluates the accuracy exclusively for unseen compositions. 3) Harmonic Mean (HM): The best harmonic mean between the seen and unseen accuracy is calculated, providing a comprehensive performance measure. 4) Area Under the Curve (AUC): This metric computes the area under the seen-unseen accuracy curve, capturing the overall performance characteristics over a wide range of operating points from $\infty$- ∞ to +$\infty$+ ∞.
Implementation Details. Please refer to the appendix.
4.2 Comparison with the State-of-the-Arts
@@
Setting. To ensure fair comparisons with existing work, we conducted a comparative analysis of PLO against several CLIP-based approaches that use the same backbone (ViT-L/14). The methods including CLIP [34], CoOp [43], Co-CGE [26], ProDA [24], PromptCompVL [37], CSP [30], all versions of DFSP [22], Troika [10], PLID [3], and GIPCOL [38]. The results are shown in Table 1222To limit the money spent on generating prompts with GPT, we only reported the results of PLO-LLM in the closed-world setting..
@@
Quantitative Analysis. It can be seen that PLO outperforms nearly all competitors by clear margins in both closed-world and open-world settings across three prevalent benchmarks. In the closed-world setting: 1) PLO-VLM surpasses the leading competitor, Troika, by 0.1%, 0.3%, and 2.1% in AUC (the core metric) on the MIT-States, UT-Zappos, and C-GQA datasets, respectively. 2) PLO-VLM and PLO-LLM achieve dominant results on C-GQA (the most challenging dataset), with HM scores of 32.2% and 31.2%, compared to Troika’s 29.4%. In the open-world setting2, PLO-VLM attains the highest AUC across all three datasets, with scores of 7.4%, 33.1%, and 3.9%. These numerical results substantiate our motivation of empowering models with the observing capabilities to understand visual compositions in a simple to complex manner, rather than modeling each composition separately.
Method C-GQA Cross-Domain
S U HM AUC
DFSP [22] 49.2 28.4 26.6 11.7
PLO-VLM 50.8 28.7 28.7 12.5
PLO-LLM 52.5 32.1 32.9 15.1
Table 2: The cross-domain evaluation on C-GQA dataset.
@@
49.6 52.7 39.0 22.2
Table 3: Ablation study on each component in PLO-VLM. PEFT: It adopts the parameter efficient fine-tuning in the image encoder. DO: It utilizes the dynamic observation strategy.
@@
Dynamic 49.6 52.7 39.0 22.2
Table 4: Ablation study on observation order in PLO-VLM.
Figure 3: Ablation study on the different number of observation prompts in PLO-LLM.
Figure 4: Top-$K$ predictions on for randomly selected cases from MIT-States. The top and bottom rows show the results of open-world settings, respectively. Correct and incorrect predictions are highlighted in green and red, respectively.
Figure 5: Distribution of samples where state or object is first observed in PLO-VLM on the test set of MIT-States.
Figure 6: Top-1 predictions of PLO-VLM with and without dynamically observing under the open-world setting on MIT-States dataset. Pre-observing primitives are highlighted in blue.
4.3 Cross-Domain Evaluation
Setting. For a more comprehensive evaluation, we trained DFSP(i2t) [22], PLO-VLM and PLO-LLM on the MIT-States dataset and tested cross-domain performance on the C-GQA dataset, selecting categories that correspond to the states and objects in MIT-States for a consistent assessment.
Results. As shown in Table 2, PLO-LLM exhibits superior performance on C-GQA dataset, particularly in the accuracy of unseen categories (32.1% vs. 28.7%) and consistently outperforms in HM and AUC scores. This enhancement stems from the GPT-generated sequence of observations in PLO-LLM, which operates independently of the visual modality, thereby bolstering cross-domain robustness.
Backbone Method MIT-States
S U HM AUC
ViT-B/32 DFSP [22] 36.7 43.4 29.4 13.2
PLO-VLM 41.1 44.2 31.3 14.8
PLO-LLM 44.2 47.9 34.3 17.4
ViT-B/16 DFSP [22] 39.6 46.5 31.5 15.1
PLO-VLM 43.9 46.6 34.1 17.0
PLO-LLM 41.2 44.7 31.4 14.8
ViT-L/14 DFSP [22] 46.9 52.0 37.3 20.6
PLO-VLM 49.6 52.7 39.0 22.2
PLO-LLM 49.6 53.2 39.0 21.9
Table 5: Ablation study on different CLIP backbones.
4.4 Ablation Study
@@
In this section, we conducted ablation studies to analyze the effectiveness of each component in PLO-VLM, the impact of the observation order in PLO-VLM, the number of observation prompts in PLO-LLM, and the network architectures of the based CLIP. All experiments were conducted on the MIT-States dataset under the closed-world setting.
@@
Effectiveness of Each Component in PLO-VLM. The contribution of Parameter Efficient Fine-Tuning (PEFT)1 and Dynamic Observation (DO) strategies to PLO-VLM was assessed, as summarized in Table 3. Our analysis can yield the following insights: 1) PEFT (Fine-Tuning in $En_{v}$): Incorporating PEFT into the image encoder $En_{v}$ marginally improves HM metric, with an increase from 36.9% to 38.1%. This improvement underscores the subtle yet impactful role of fine-tuning the encoder’s lightweight layers. Extra experimental results of PEFT are left in the appendix. 2) DO: Implementing DO in isolation enhances the model’s ability to recognize both seen and unseen compositions, as reflected in a boost of the seen metric from 47.6% to 48.5% and unseen metric from 51.2% to 52.0%. This highlights the critical impact of dynamically determining observations in understanding visual compositions. 3) Synergistic Impact of PEFT and DO: When PEFT and DO are synergistically applied, they collectively achieve the highest AUC of 22.2%. This composite application underlines the effectiveness of integrating both strategies, leading to optimal model performance in recognizing compositions.
@@
Observation Order. We explored the influence of different observation orders in PLO-VLM by using various strategies for determining the first observation: predicting the state primitive first, predicting the object primitive first, and dynamically deciding the observation order. From Table 4, we can observe: 1) The multi-step observation leads to significant performance gains against the baseline across all the metrics. 2) Dynamically deciding the observation order based on the input image yields the highest overall performance. These findings demonstrated the effectiveness of adaptively choosing the most informative observation step based on the content of the input image, which is of importance due to the conditioned variance nature of CZSL.
@@
Figure 5 reveals the dynamic observation strategy of PLO-VLM, which selects states or objects first based on their visual saliency in a composition [^3] More visualization results are left in the appendix.. The “cooked pasta” often undergoes object-first observations owing to its distinct visuals, whereas “broken bottle” typically requires state-first recognition due to the impactful visual change presented by the state “broken”. PLO-VLM’s context-aware approach mirrors human cognition in compositional zero-shot learning, leading to more precise recognition. Furthermore, we visualized top-1 predictions of PLO-VLM with and without dynamically observing in Figure 6. In a sense, the pre-observed primitive can represent a more prominent feature in the image (e.g., “mashed object” and “burnt object”). By virtue of pre-observed primitive, PLO-VLM consistently achieved accurate predictions.
@@
Number of Observation Prompts. We investigated the impact of the number of observation prompts in PLO-LLM. We varied the number of prompts for each composition category and evaluated the model’s performance accordingly. The results, depicted in Figure 3, reveal a nuanced relationship between the number of prompts and the model’s accuracy. As expected, increasing the number of observation prompts can generally lead to improved performance.
Effect of Network Architectures. We further examined the influence of replacing CLIP backbones in our PLO-VLM and PLO-LLM. All results are reported in Table 5. Those consistent performance gains compared to the previous SOTA reaffirmed the efficacy and robustness of the proposed methodology. We chose ViT-L/14 by default.
Qualitative Results. We provided top-$K$ predictions of PLO-VLM in the open-world settings in Figure 4. Our progressive observation strategy enables a thorough grasp of compositions, which is particularly effective in recognizing unseen compositions by bridging the gap between base and novel categories. Notably, even when our model’s top-1 prediction is not exactly matched, it still accurately predicts the state primitive, and the object primitive also appears to be reasonable (e.g., muddy stream). These results underscore PLO’s effectiveness in capturing pertinent visual cues and enabling comprehensive composition understanding.
5 Conclusion
@@
In this paper, we presented PLO, a novel solution for addressing the challenges of conditioned variance in CZSL. Unlike existing methods modeling each composition separately, PLO effectively captures the interactions between state and object primitives by dynamically determining the observation order, leading to a holistic understanding of visual concepts for each composition. The two variants, PLO-VLM and PLO-LLM, harness the power of VLMs and LLMs to refine image representations in a simple to complex manner. Experimental results on three gold-standard CZSL benchmarks demonstrated their superiority over existing frameworks. PLO opens a new avenue for CZSL from the perspective of endowing models with observing capabilities, and we hope it will pave the way for future research.
Supplementary Material
@@
This supplementary document is organized as follows:
@@
  • Implementation details referenced in Sec. 4.1 are provided in Sec. C.
  • Additional quantitative results are presented in Sec. D.
  • Additional qualitative results are displayed in Sec. E.
  • Potential negative societal impacts are discussed in Sec. F.
  • Limitations are considered in Sec. G.
A The procedures of Training and Inference
@@
B Observation Prompt Generation
We present the prompt designed to generate a series of observation prompts that fed into the text encoder of CLIP within the framework of PLO-LLM. The prompt’s structure is as follows:
@@ B Observation Prompt Generation
   - a photo of yellow, mushy substance
   - a photo of a fruit that has been mashed into a paste
   - a photo of a soft and creamy mixture made from bananas
   - a photo of mashed banana
@@
Data: Training data $DS=\{(I,c)\}$
Result: Optimal parameters for PLO-VLM
1 Initialization: Initialize weights;
// Training Process
2 while not converged do
3       Sample a batch from $DS$ with images $\{I_{k}\}_{k=1}^{n}$ and their corresponding labels $\{c_{k}\}_{k=1}^{n}$;
4       Extract $\mathbf{v}$, $\mathbf{t}_{s}$, $\mathbf{t}_{o}$, and $\mathbf{t}_{c}$;
5       Use the pre-observing detector $\mathcal{F}^{pre}_{obs}(\mathbf{v},\mathbf{t}_{s},\mathbf{t}_{o})$ to determine observation order;
6       Compute $\mathcal{L}_{obs}$; if observing object first then
7             Obtain $\mathbf{t}^{pre}_{obs}\leftarrow\mathbf{t}_{o}$, $\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{s}$ ;
8             Compute refined $\widetilde{\mathbf{v}}$ via CA, based on $\mathbf{t}^{pre}_{obs}$;
9             Compute $p(s|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs})$ ;
10             Compute loss for state: $\mathcal{L}_{s}$;
11
12      else
13             Obtain $\mathbf{t}^{pre}_{obs}\leftarrow\mathbf{t}_{s}$, $\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{o}$ ;
14             Compute refined $\widetilde{\mathbf{v}}$ via CA, based on $\mathbf{t}^{pre}_{obs}$;
15             Compute $p(o|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs})$ ;
16             Compute loss for object: $\mathcal{L}_{o}$;
17
18      Compute $p(c|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}_{c})$ and $\mathcal{L}_{c}$;
19       Calculate total loss: $\mathcal{L}_{PLO}^{VLM}=\mathcal{L}_{obs}+\mathcal{L}_{s}+\mathcal{L}_{o}+\mathcal{L}_{c}$;
20       Update weights using $\mathcal{L}_{PLO}^{VLM}$;
21
// Inference Process
22 for each test image I do
23       Compute $p(c|I)$;
24       Predict label: $\hat{y}_{c}=\text{argmax }p(c|I)$;
25
@@ 20       Update weights using $\mathcal{L}_{PLO}^{VLM}$;
Data: Training data $DS=\{(I,c)\}$
Result: Optimal parameters for PLO-LLM
1 Initialization: Initialize weights;
2 Observation Prompts Generation:
3 for each composition category $c\in C$ do
4       Generate observation prompts $\mathcal{P}_{c}$ using LLM;
5
// Training Process
6 while not converged do
7       Sample a batch from $DS$ with images $\{I_{k}\}_{k=1}^{n}$ and their corresponding labels $\{c_{k}\}_{k=1}^{n}$;
8       for each observation step $i$ do
9             Extract $\mathbf{v}^{(i)}$ and $\mathbf{t}^{(i)}_{obs}$;
10             Compute refined $\widetilde{\mathbf{v}}^{(i)}$ via CA and update $\mathbf{v}^{(i+1)}$;
11             Compute probability: $p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)})$;
12
13      Calculate total loss: $\mathcal{L}_{PLO}^{LLM}=\mathcal{L}_{step}+\mathcal{L}_{c}$;
14       Update weights using $\mathcal{L}_{PLO}^{LLM}$;
15
// Inference Process
16 for each test image I do
17       Compute $p^{Soft}(c|I)$;
18       for each observation step $i$ do
19             Update $p^{Hard}(c|I)$ using hard prompts;
20
21      Predict label: $\hat{y}_{c}=\text{argmax }p(c|I)$;
22
@@ 14       Update weights using $\mathcal{L}_{PLO}^{LLM}$;
  • Setting: The setting text (i.e., “In compositional classification...”) establishes a specific context and roles for the Large Language Models (LLMs) to operate within.
@@
  • Example: The provided instance (i.e., the example of “mashed banana”) functions as a guiding paradigm for the model to generate analogous output in the manner of in-context learning [4, 21].
  • Question: The question (i.e., “How can you identify...”) instructs the model to devise observation prompts tailored to the specific composition category under consideration.
In addition, this process adheres to the Chain-of-Thought concept [16, 41], generating observations step by step from easy to hard through the utilization of a special prompt “Let’s observe it step by step!”.
Figure 7: Distribution of samples where state or object is first observed in PLO-VLM on the test set of MIT-States.
Figure 8: Top-$K$ predictions on for randomly selected cases from three datasets. The top and bottom rows show the results of closed-world settings, respectively. Correct and incorrect predictions are highlighted in green and red, respectively.
Figure 9: Top-1 predictions of PLO-LLM with and without multi-step observations under the closed-world setting on the MIT-States dataset. The corresponding generated observation prompts are presented to the right of each image. Correct and incorrect predictions are highlighted in green and red, respectively.
C Implementation Details
Our PLO models were trained and evaluated on one NVIDIA A100 GPU using PyTorch [32]. The GPT-3.5-turbo, a variant of the GPT model, known for its impressive performance, was employed as the LLM. For the CLIP, we utilized OpenAI’s resources, opting for the Vision Transformer with a base configuration of ViT-L/14. In PLO-VLM, we assigned the weight factors for $\mathcal{L}_{obs}$, $\mathcal{L}_{s}$, $\mathcal{L}_{o}$, and $\mathcal{L}_{c}$ to 1.0, 0.01, 0.01, and 1.0, respectively, as also presented in [22]. In PLO-LLM, the default number of observation prompts was set to 4. Further, fusion weights during the aggregation of probabilities for $p^{Soft}(c|I)$ and $p^{Hard}(c|I)$ were set to 0.7 and 0.3, respectively. Moreover, the weight factors assigned to $\mathcal{L}_{step}$ and $\mathcal{L}_{c}$ were both set to 1.0. Following [22], we used the Adam optimizer for 20 epochs on all datasets, with a learning rate of 2.5e-4 and a batch size of 64. In the open-world evaluation, we adhered to the post-training calibration method [30] to filter out compositions deemed infeasible. Our code will be released.

### v2

@@ 4.1. Experiment Settings
Datasets. We evaluated performance on three challenging benchmarks: MIT-States (Isola et al., 2015): It comprises 53,753 natural images, with 115 states and 245 objects. In the closed-world setting, the search space includes 1,262 seen compositions and 300/400 unseen compositions for validation/testing. 2) UT-Zappos (Naeem et al., 2021): It consists of 50,025 images of shoes, with 16 states and 12 objects. In the closed-world experiments, we considered 83 seen compositions and 15/18 (validation/test) unseen compositions following the constraints defined in (Purushwalkam et al., 2019). 3) C-GQA (Yu and Grauman, 2014): It contains 453 states and 870 objects, comprising a total of 39,298 images. The dataset is divided into 5,592 seen compositions for training, and 1,040/923 unseen compositions for validation/testing, respectively. In the open-world settings, these datasets contain 28,175, 192, and 278,362 compositions, respectively.
@@
Metrics. We followed the established CZSL evaluation protocol (Mancini et al., 2021) and assessed all results using four metrics in both closed-world and open-world scenarios: 1) Seen (S): measures the accuracy specifically for seen compositions. 2) Unseen (U): evaluates the accuracy exclusively for unseen compositions. 3) Harmonic Mean (HM): denotes the best harmonic mean between the seen and unseen accuracy, providing a comprehensive performance measure. 4) Area Under the Curve (AUC): computes the area under the seen-unseen accuracy curve, capturing the overall performance characteristics over a wide range of operating points from -$\infty$ to $\infty$.
Implementation Details. Our PLO models were trained and evaluated on one NVIDIA A100 GPU using PyTorch (Paszke et al., 2019). The GPT-3.5-turbo, a variant of the GPT model, known for its impressive performance, was employed as the LLM. For the CLIP, we utilized OpenAI’s resources, opting for the Vision Transformer with a base configuration of ViT-L/14. In PLO-VLM, we assigned the weight factors for $\mathcal{L}_{obs}$, $\mathcal{L}_{s}$, $\mathcal{L}_{o}$, and $\mathcal{L}_{c}$ to 1.0, 0.01, 0.01, and 1.0, respectively, as also presented in (Lu et al., 2023a). In PLO-LLM, the default number of observation cues was set to 4. Further, fusion weights during the aggregation of probabilities for $p^{Soft}(c|I)$ and $p^{Hard}(c|I)$ were set to 0.7 and 0.3, respectively. Moreover, the weight factors assigned to $\mathcal{L}_{step}$ and $\mathcal{L}_{c}$ were both set to 1.0. Following (Lu et al., 2023a), we used the Adam optimizer for 20 epochs on all datasets, with a learning rate of 2.5e-4 and a batch size of 64. In the open-world evaluation, we adhered to the post-training calibration method (Nayak et al., 2023) to filter out compositions deemed infeasible.
4.2. Comparison with the State-of-the-Arts
@@
Setting. To ensure fair comparisons with existing work, we conducted a comparative analysis of PLO against several CLIP-based approaches that use the same backbone (ViT-L/14). The methods including CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022b), Co-CGE (Mancini et al., 2022), ProDA (Lu et al., 2022), CSP (Nayak et al., 2023), all versions of DFSP (Lu et al., 2023a), GIPCOL (Xu et al., 2023), CAILA (Zheng et al., 2024), Troika (Huang et al., 2023), CDS-CZSL (Li et al., 2024) and PLID (Bao et al., 2024).
Quantitative Analysis. The results are shown in Table 1.
@@
It can be seen that PLO outperforms nearly all competitors by clear margins in both closed-world and open-world settings across three prevalent benchmarks. In the closed-world setting: 1) PLO-VLM surpasses the leading competitor, Troika (with the official adapter (Houlsby et al., 2019)), by 0.2%, 0.3%, and 2.1% in AUC (the core metric) on the MIT-States, UT-Zappos, and C-GQA datasets, respectively. PLO-VLM outperforms the existing SOTA methods on almost all metrics. 2) PLO-VLM and PLO-LLM achieve dominant results on C-GQA (the most challenging dataset), with HM scores of 32.2% and 31.2%, compared to Troika’s 29.4%. In the open-world setting: 1) PLO-VLM attains the highest AUC across two datasets, with scores of 7.8%, 33.1%, and 3.9%. The performance gains may be that open-world is a more difficult setting with no constraint on the test time search space (Mancini et al., 2021). Our PLO can achieve better performance by implicitly contracting the search space by first observing the salient features. 2) PLO-LLM also exhibits strong performance, particularly excelling in the UT-Zappos dataset (48.8% on HM metric). Notably, since the visual image properties (e.g., resolution, clarity, and consistency) of the unseen portion in the same dataset align with those of the seen data, PLO-VLM marginally outperforms PLO-LLM on these samples. However, for cross-domain data (cf., Appendix), PLO-LLM significantly surpasses PLO-VLM. These numerical results substantiate our motivation of empowering models with the observing capabilities to understand visual compositions in a simple-to-complex manner, rather than modeling each composition separately.
@@
49.7 52.8 39.4 22.3
(a) Ablation study on each component.
@@
Dynamic 49.7 52.8 39.4 22.3
(b) Ablation study on observation order.
Table 2. Ablation study on PLO-VLM.
4.3. Ablation Study
@@
In this section, we conducted ablation studies to analyze the effectiveness of each component in PLO-VLM, the impact of the observation order in PLO-VLM, the number of observation cues in PLO-LLM, and the network architectures of the based CLIP. All experiments were conducted on the MIT-States dataset under the closed-world setting [^1] More ablation study results are left in the Appendix..
@@
Key Components in PLO-VLM. The contribution of Parameter Efficient Fine-Tuning (PEFT) and Dynamic Observation (DO) strategies to PLO-VLM was assessed, as summarized in Table 2(a). Our analysis can yield the following insights: 1) PEFT (Fine-Tuning in $En_{v}$): Incorporating PEFT into the image encoder $En_{v}$ marginally improves HM metric, with an increase from 36.9% to 38.1%. This improvement underscores the subtle yet impactful role of fine-tuning the encoder’s lightweight layers. More results of PEFT are left in the Appendix. 2) DO: Implementing DO in isolation enhances the model’s ability to recognize both seen and unseen compositions, as reflected in a boost of the seen metric from 47.6% to 48.5% and unseen metric from 51.2% to 52.0%. Even when compared to the SOTA DFSP method without PEFT, our DO strategy achieves a certain improvement (21.3% vs. 20.8% on AUC metric). This highlights the critical impact of dynamically determining observations in understanding visual compositions. 3) Synergistic Impact of PEFT and DO: When PEFT and DO are synergistically applied, they collectively achieve the highest AUC of 22.3%. This composite application underlines the effectiveness of integrating both strategies, leading to optimal model performance in recognizing compositions.
Figure 3. Distribution of samples where state or object is first observed in PLO-VLM on MIT-States test set.
Figure 4. Ablation study on the different number of observation cues in PLO-LLM.
@@
Observation Order. We explored the influence of different observation orders in PLO-VLM by using various strategies for determining the first observation: predicting the state primitive first, predicting the object primitive first, and dynamically deciding the observation order. From Table 2(b), we can observe: 1) The multi-step observation leads to significant performance gains against the baseline across all the metrics. 2) Dynamically deciding the observation order based on the input image yields the highest overall performance. These findings verified the effectiveness of adaptively choosing the most informative observation step based on the content of the input image, which is of importance due to the conditioned variance nature of CZSL.
@@
Figure 3 reveals the dynamic observation strategy of PLO-VLM, which selects states or objects first based on their visual saliency in a composition. Note that in the stacked bar, blue represents the samples of “state first”, green represents “object first”, and the green bar DOES NOT cover blue samples. As seen, the “cooked pasta” often undergoes object-first observations owing to its distinct visuals, whereas “broken bottle” typically requires state-first recognition due to the impactful visual change presented by the state “broken”. PLO-VLM’s context-aware approach mirrors human cognition, leading to more precise recognition.
@@
Number of Observation Cues. We investigated the impact of the number of observation cues in PLO-LLM. We varied the number of cues for each composition category and evaluated the model’s performance accordingly. The results in Figure 4 reveal a nuanced relationship between the number of cues and the model’s accuracy. As expected, increasing the number of observed cues can generally lead to improved performance, and performance is saturated when the number is 4. Thus, the number of cues is set to 4 as default.
Figure 5. Top-1 predictions of PLO-VLM with and w/o dynamically observing under open-world setting on MIT-States. Pre-observing primitives are highlighted in blue.
4.4. Visualization
In this section, the validity of PLO is further illustrated by the prediction results of some samples and the cross-attention map visualization of each observation step.
Top-1 Predictions. We visualized top-1 predictions of PLO-VLM with and without dynamically observing in Figure 5. In a sense, the pre-observed primitive can represent a more prominent feature in the image (e.g., “mashed object” and “burnt object”). By virtue of pre-observed primitive, PLO-VLM consistently achieved accurate predictions. In Figure 6, we presented the top-1 predictions obtained by PLO-LLM, in cases where observations (obs) were utilized or not. Concurrently, we displayed the sequence of step-by-step observation cues on the right side of each image. With our meticulously designed prompt, the LLM effectively generated a progression of observation cues that transitioned from easy to hard. Through the aid of these well-organized observation cues, we empowered the model with observing capabilities, leading to better holistic understanding and considerable performance gains.
Figure 6. Top-1 predictions of PLO-LLM with and w/o multi-step observations under the closed-world setting on the MIT-States dataset. Corresponding observation cues are on the right of each image. Correct and incorrect predictions are in green and red, respectively.
Figure 7. Visualization of cross-attention maps based on different primitives in PLO-VLM.
Figure 8. Visualization of cross-attention maps at different observation steps in PLO-LLM.
Cross-Attention Map. We displayed the cross-attention maps in Eq. (4) of different primitives in PLO-VLM (cf., Figure 7) and of different steps in PLO-LLM (cf., Figure 8) to explain our reasons for the performance improvement.
PLO-VLM. 1) State First: When our model detects that the state features are more prominent, e.g., the mushy texture common to various “pureed” samples. PLO-VLM uses cross-attention to focus on the color and texture in the “pureed” areas (cf., Figure 7(a)). This allows it to correctly identify the composition as “pureed berry”. Methods that rely solely on object features might misclassify this image as “fresh berry” (cf., Figure 7(b)). 2) Object First: Conversely, when object features are more discernable, such as the crescent shape of a banana, the model also applies cross-attention to look for signs of ripeness, like brown spots, to determine the composition as “ripe banana” (cf., Figure 7(c)). In this scenario, focusing on less prominent state features may lead to misclassification (cf., Figure 7(d)). By recognizing the most salient state/object automatically, our method bypasses the drawbacks of fixed order strategies, leading to more accurate recognition of compositions.
PLO-LLM. It can be seen that the regions of interest gradually coalesce as the steps progress. Taking the composition “caramelized apple” (top row) as an example, with the ordered text cues (i.e., “golden-brown”, “with sugar”, and “caramel syrup”), our model gradually attends to the distinctive regions of “caramelized apple”, suggesting the explainability and effectiveness.
5. Conclusion
@@
In this paper, we presented PLO, a novel solution for addressing the challenges of conditioned variance in CZSL. Unlike existing methods modeling each composition separately, PLO effectively captures the interactions between state and object primitives as well as graduated descriptions by automatically determining the observation cue order, leading to a holistic understanding of visual concepts for each composition. The two variants, PLO-VLM and PLO-LLM, harness the power of VLMs and LLMs to refine image representations in a simple-to-complex manner. Experimental results on three gold-standard CZSL benchmarks demonstrated their superiority over existing frameworks. PLO opens a new avenue for CZSL from the perspective of endowing models with observing capabilities, and we hope it will pave the way for future research.
6. Acknowledgments
This work was supported by the National Key Research and Development Program of China (2024YFB3312900), Zhejiang Provincial Natural Science Foundation of China (No. LD25F020001), and the Fundamental Research Funds for the Central Universities (226-2025-00057). Long Chen was supported by the Hong Kong SAR RGC Early Career Scheme (26208924), the National Natural Science Foundation of China Young Scholar Fund (62402408), and the HKUST Sports Science and Technology Research Grant (SSTRG24EG04). This research was partially conducted by ACCESS – AI Chip Center for Emerging Smart Systems, supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government.
@@ 6. Acknowledgments
This appendix is organized as follows:
@@
  • Additional quantitative results are presented in Sec. C.
  • Additional qualitative results are displayed in Sec. D.
  • Discussion between two PLO variants is shown in Sec. E.
  • Limitations are considered in Sec. F.
Appendix A The procedures of Training and Inference
@@
Appendix B Observation Cue Generation
We present the prompt designed to generate a series of observation cues that fed into the text encoder of CLIP within the framework of PLO-LLM. The prompt’s structure is as follows:
@@ Appendix B Observation Cue Generation
- a photo of yellow, mushy substance
- a photo of a fruit that has been mashed into a paste
- a photo of a soft and creamy mixture made from bananas
- a photo of mashed banana
@@
1:Input: Training data $DS=\{(I,c)\}$
2:Output: Optimal parameters for PLO-VLM
3:Initialization: Initialize weights
4:// Training Process
5:while not converged do
6:  Sample a batch from DSDS with images $\{I_{k}\}_{k=1}^{n}$ and their corresponding labels $\{c_{k}\}_{k=1}^{n}$
7:  Extract $\mathbf{v}$, $\mathbf{t}_{s}$, $\mathbf{t}_{o}$, and $\mathbf{t}_{c}$
8:  Use the pre-observing detector $\mathcal{F}^{pre}_{obs}(\mathbf{v},\mathbf{t}_{s},\mathbf{t}_{o})$ to determine observation order
9:  Compute $\mathcal{L}_{obs}$
10:  if observing object first then
11:   Obtain $\mathbf{t}^{pre}_{obs}\leftarrow\mathbf{t}_{o}$, $\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{s}$
12:   Compute refined $\widetilde{\mathbf{v}}$ via CA, based on $\mathbf{t}^{pre}_{obs}$
13:   Compute $p(s|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs})$
14:   Compute loss for state: $\mathcal{L}_{s}$
15:  else
16:   Obtain $\mathbf{t}^{pre}_{obs}\leftarrow\mathbf{t}_{s}$, $\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{o}$
17:   Compute the refined image representation $\widetilde{\mathbf{v}}$ via CA, based on $\mathbf{t}^{pre}_{obs}$
18:   Compute $p(o|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs})$
19:   Compute loss for object: $\mathcal{L}_{o}$
20:  end if
21:  Compute $p(c|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}_{c})$ and $\mathcal{L}_{c}$
22:  Calculate total loss: $\mathcal{L}_{PLO}^{VLM}=\mathcal{L}_{obs}+\mathcal{L}_{s}+\mathcal{L}_{o}+\mathcal{L}_{c}$
23:  Update weights using $\mathcal{L}_{PLO}^{VLM}$
24:end while
25:// Inference Process
26:for each test image II do
27:  Compute $p(c|I)$
28:  Predict label: $\hat{y}_{c}=\text{argmax }p(c|I)$
29:end for
@@
1:Input: Training data $DS=\{(I,c)\}$
2:Output: Optimal parameters for PLO-LLM
3:Initialization: Initialize weights
4:Observation Cues Generation:
5:for each composition category $c\in C$ do
6:  Generate observation cues $\mathcal{P}_{c}$ using LLM
7:end for
8:// Training Process
9:while not converged do
10:  Sample a batch from DSDS with images $\{I_{k}\}_{k=1}^{n}$ and their corresponding labels $\{c_{k}\}_{k=1}^{n}$
11:  for each observation step $i$ do
12:   Extract $\mathbf{v}^{(i)}$ and $\mathbf{t}^{(i)}_{obs}$
13:   Compute the refined image representation $\widetilde{\mathbf{v}}^{(i)}$ via CA and update $\mathbf{v}^{(i+1)}$
14:   Compute probability: $p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)})$
15:  end for
16:  Calculate total loss: $\mathcal{L}_{PLO}^{LLM}=\mathcal{L}_{step}+\mathcal{L}_{c}$
17:  Update weights using $\mathcal{L}_{PLO}^{LLM}$
18:end while
19:// Inference Process
20:for each test image II do
21:  Compute $p^{Soft}(c|I)$
22:  for each observation step $i$ do
23:   Update $p^{Hard}(c|I)$ using hard prompts
24:  end for
25:  Predict label: $\hat{y}_{c}=\text{argmax }p(c|I)$
26:end for
@@
  • Setting: The configuration setting text (i.e., “In compositional classification...”) establishes a specific context and roles for the Large Language Models (LLMs) to operate within.
@@
  • Example: The provided instance (i.e., the example of “mashed banana”) functions as a guiding paradigm for the model to generate analogous output in the manner of in-context learning (Brown et al., 2020; Liu et al., 2021; Shao et al., 2025a).
  • Question: The question (i.e., “How can you identify...”) instructs the model to devise observation cues tailored to the specific composition category under consideration.
  Method C-GQA MIT-States
S U HM AUC S U HM AUC
DFSP (Lu et al., 2023a) 49.2 28.4 26.6 11.7 45.7 42.8 33.9 16.8
PLO-VLM 50.8 28.7 28.7 12.5 46.2 42.4 35.3 17.2
PLO-LLM 52.5 32.1 32.9 15.1 50.5 43.6 37.9 19.4
Table 3. Cross-domain evaluation on C-GQA and MIT-States.
  Strategy MIT-States
S U HM AUC
CLS-DES (Menon and Vondrick, 2023) 47.0 51.8 37.3 20.3
PLO-LLM 49.6 53.2 39.0 21.9
Table 4. Comparison of LLM-based strategies on the MIT-States dataset.